OpenStudy (anonymous):

please integrate this #1.) x^3 squareroot of (a^2+x^2) dx #2.)dx/x(x^2-a)^2 #3.)z^5 dz/(z^2-9^2)

OpenStudy (anonymous):

For number 1, let u=x^3 so we can get rid of the x^3 to use the arcsin formula. The arcsin formula is integral of du/sqrt of a^2-u^2= arcsin u/a + c.

OpenStudy (anonymous):

can i ask is the a here has a value?

OpenStudy (anonymous):

the first one you should sub u=a^2+x^2

OpenStudy (anonymous):

the other guy is wrong :|

OpenStudy (anonymous):

then? can u plaease help me out?

OpenStudy (anonymous):

\[\int\limits x^3 \sqrt{a^2 +x^2} dx \] if its this

OpenStudy (anonymous):

u=a^2 +x^2 du/dx= 2x so x dx = du /2

OpenStudy (anonymous):

rearranging the substitution , x^2 = u-a^2

OpenStudy (anonymous):

rewrite the integral \[\int\limits x \times x^2 \times \sqrt{a^2+x^2} dx = \int\limits x^2 \times \sqrt{a^2+x^2} \times (x dx ) \]

OpenStudy (anonymous):

is this integral by parts?

OpenStudy (anonymous):

\[= \int\limits (u-a^2) \times \sqrt{u} \times \frac{du}{2} = \frac{1}{2} \int\limits (u^{\frac{3}{2}} - a^2 u^{\frac{1}{2}} )du \]

OpenStudy (anonymous):

no by substitution

OpenStudy (anonymous):

its easy from then ^

OpenStudy (anonymous):

a^2 is just a constant

OpenStudy (anonymous):

thank u. wat about the number 2 and 3?

OpenStudy (anonymous):

3 is by partial fractions

OpenStudy (anonymous):

wait, thats wrong

OpenStudy (anonymous):

you need to divide first , because the degree of the numerator is larger than the denominator

OpenStudy (anonymous):

its z^5 /(z^2-a^2)dz

OpenStudy (anonymous):

the second one is partial fractions as well \[= \int\limits \frac{dx}{ x (x-\sqrt{a})^2 (x+\sqrt{a})^2} = \int\limits \frac{A}{x} + \frac{B}{x-\sqrt{a}} +\frac{C}{ (x-\sqrt{a})^2} + \frac{D}{(x+\sqrt{a})} + \frac{E}{(x+\sqrt{a})^2}\]